### v5

@@ Abstract
Penetration testing allows organizations to identify and remediate vulnerabilities in their systems. It is an essential part of active software security testing as it bolsters defense mechanisms against cyberattacks. One recent advancement in the realm of penetration testing is the utilization of Language Models (LLMs). We explore the intersection of LLMs and penetration testing to gain insight into their capabilities and challenges, esp. in the context of privilege escalation.
We introduce a fully automated privilege-escalation tool designed for evaluating the efficacy of LLMs for ethical hacking, executing benchmarks using multiple LLMs, and investigating their respective results.
A localized benchmark allows us to evaluate the effectiveness of privilege-escalation techniques performed by both humans and automated tooling. To address this gap, we developed a comprehensive benchmark for Linux privilege escalation. It provides a standardized platform to evaluate and compare the performance of human and synthetic actors, e.g., hacking scripts or automated tooling in the form of an agentic LLM system.
We analyze the impact of different context sizes, in-context learning, high-level guidance mechanisms, and memory management techniques. We discuss challenging areas for LLMs, including maintaining focus during testing, coping with errors, and finally comparing LLMs with human hackers. Our results show that GPT-4-turbo is well suited to exploit vulnerabilities (33–83% of vulnerabilities). GPT-3.5-turbo can abuse 16–50% of vulnerabilities, while local models, such as Llama3, can only exploit between 0 and 33% of the vulnerabilities.
The current version of the LLM-guided privilege-escalation prototype can be found at https://github.com/ipa-labs/hackingBuddyGPT. Our Linux Privilege Escalation benchmark can be retrieved from https://github.com/ipa-labs/benchmark-privesc-linux. Both are already in use by third parties.
@@ 1 Introduction
In the rapidly evolving field of cybersecurity, penetration testing (“pen-testing” or “hacking”) plays a pivotal role in identifying and mitigating potential vulnerabilities. A crucial subtask of pen-testing is privilege escalation, which involves exploiting a bug, design flaw, or configuration oversight in an operating system or software application to gain elevated access to resources that are normally protected from an application or user [^1] https://en.wikipedia.org/wiki/Privilege_escalation. The ability to escalate privileges provides a malicious actor with increased access, potentially leading to more significant breaches or system damage. Therefore, understanding and improving the performance of tools used for this task is highly relevant and impacts real-life security.
@@
In this paper, we focus on investigating the performance of Large Language Models (LLMs) in the context of penetration testing, specifically for Linux privilege escalation. LLMs have shown remarkable abilities to emulate human behavior that can be used to automate and improve various tasks in pen-testing (Deng et al., 2023; Happe and Cito, 2023a). However, there is currently no understanding on how these models perform in common privilege escalation scenarios.
@@
To address this gap, we performed an empirical analysis of multiple LLMs using an open-source Linux privilege-escalation benchmark, providing insight into LLMs’ strengths and weaknesses in the context of these attacks. We release a platform to evaluate and compare the performance of different LLMs in a controlled manner. By understanding the performance of these models in the critical task of privilege escalation, we can guide future research efforts towards higher effectiveness and reliability for LLM-guided penetration testing.
Contributions.
@@
This work arose from the question “What is the efficacy of LLMs for Linux Privilege-Escalation Attacks”? To answer it, we implemented an LLM-driven exploitation tool designed for rapid prototyping, and identified properties of LLM-based penetration testing through empirical analysis of benchmark runs performed. This approach results in the following contributions:
@@
  • a publicly available Linux Privilege Escalation benchmark set that can be run on local premises due to the safety- and security-critical nature of this benchmark (Section 4).
@@
  • an fully-automated LLM-driven Linux privilege escalation prototype (Section 3 Methodology). We release the prototype including all prompts on github as open-source.
@@
While these detail our high-level goals, some of the questions we’ve answered within this work are:
  • We show that autonomous penetration-testing using a single LLM-callout control loop is a resource-efficient alternative to commonly used multi-LLM agent based systems. Compared to state-of-the-art autonomous multi-LLM solutions which achieve 7–20% hacking success rates, we achieve up to 66% in unguided scenarios. This is comparable to a human baseline of 75%.
  • When adding high-level guidance, similar to the guidance provided by human penetration testers or the Planner module of pentestGPT-inspired architectures, the success rate rises to 83%. This is comparable to the success rate of our human baseline of 91%.
  • The state of the art offers conflicting guidance about the efficacy of adding background hacking information as well as on the impact of using larger execution histories. This work investigates these questions empirically, indicating that background hacking information has a smaller impact when using state-of-the-art models such as GPT-4 while increasing the stored history increased our success rates significantly.
@@ 2 Background and Related Work
The background section focuses on the two distinct areas that this work integrates: LLMs and penetration testing, also known as “hacking”.
@@ 2.1 Large-Language Models
There is an ongoing discussion about the minimum viable model parameter sizes. On the one hand, proponents claim that emergent features arise only with larger model sizes (Kosinski, 2023; Bubeck et al., 2023; Wei et al., 2022); on the other hand, proponents claim that smaller models can achieve domain-specific tasks with reduced costs for both training and execution (Bender et al., 2021). Smaller models are feasible to run locally. This is important for agent-based scenarios (Andreas, 2022; Park et al., 2023) or if privacy reasons disallow the usage of cloud-based LLMs. In early 2024 the term Small Language Models was used to denote models with parameter sizes smaller than 8 billions, one example of such a model would be Llama3-8b.
An alternative to using small language models is quantizing models with larger parameter counts. In this approach, parts of the model weights are quantized from 32bit floating points into data types of lower precision, e.g., 4 bit integers. This reduces the model’s memory requirements, and thus makes local model usage computationally feasible. There is an ongoing discussion on the relative trade-off between using smaller models of full precision and using larger quantized models (Huang et al., 2024).
Training an LLM incurs high costs. Recently, alternative approaches have tried to achieve high performance while avoiding expensive training. In-Context Learning (Dong et al., 2022; Dai et al., 2023) includes background information within the prompt and thus exchanges knowledge inherently stored within the model with external knowledge. Similarly, Chain-of-Thought prompting includes step-by-step answer examples within the context (Kojima et al., 2022). Both approaches make the context a very limited resource, even with ever increasing context-sizes of newer model generations.
Real-world tasks often must be split into smaller subtasks or steps. Multiple approaches try to emulate this through LLMs, ranging from minimal approaches such as BabyAGI (Nakajima, 2023) to Tree-of-Thoughts (Yao et al., 2023) or Task-Lists (Deng et al., 2023). Wang et al. (2023) name this approach in a contemporary paper plan-and-solve.
A combination of the mentioned topics, i.e., small viable model sizes, using context for adding information while having enough context to describe the task at hand and having task/state management for keeping track of sophisticated work, would make LLMs viable for local usage, esp. with sensitive locally-stored data.
Another problem is the missing explainabiliy of LLMs. While initial forays exist (Merullo et al., 2023), they are currently only applicable to small and out-dated LLMs. Currently, no a priori logical analysis of a LLM’s capabilities is possible, we can only perform empirical research.
LLM Benchmarks. LLM benchmarks are typically based on common sense reasoning tasks. This is sensible, as common-sense reasoning is a transferable skill well suited to many tasks, including penetration-testing. However, a recent survey by Davis (2023) shows that many existing common sense reasoning benchmarks have quality issues within their tasks. Another open question is if high scores in synthetic common-sense benchmarks translate into high scores in real-world domain-specific scenarios, which are typically not tested by LLM makers.
@@ 2.2 Penetration Testing
Penetration Testing, short pen-testing, is described by Geer and Harthorne (2002) as “the art of finding an open door”. Its goal is to find a vulnerability within the subject-under-test to falsify the hypothesis that the subject is secure. The outcome of a penetration test allows defenders to fortify their systems so that other, potentially malicious, attackers cannot abuse similar vulnerabilities (Bishop, 2007). Professionals performing those tests are typically called penetration-testers, pen-tester, or simply hackers. An additional differentiation is often performed upon the intend of the pen-tester: white hats perform ethical research to improve the field of software security while black-hats are malicious and work for monetary or political gain.
@@
Shah and Mehtre (2015) further elaborate on the nature of penetration testing and differentiate between Vulnerability Assessments and Penetration Testing. The goal of the former is to identify as many possible vulnerabilities within the subject-under-test as possible, while the latter emulates an attacker that tries to actively exploit a found vulnerability. As Penetration Testing can lead to system instabilities and data loss, automated tooling often focus upon well-known Vulnerability Assessment, not exploitation (Shebli and Beheshti, 2018). Tooling such as nmap [^2] https://nmap.org/, OpenVAS [^3] https://www.openvas.org/, PortSwigger BURP [^4] https://portswigger.net/burp or ZAP [^5] https://www.zaproxy.org/ often utilize rule-based detection systems as well as databases of known vulnerable software versions and use more-aggressive techniques such as fuzzing only upon explicit user interaction and as measures of last resort.
@@
Only little empirical research into how penetration testers perform their work, and the potential problems therein, has been performed. Happe and Cito (2023b) performed an interview study with professional penetration testers. One of their key findings was that security researchers and security practitioners (penetration testers) differ in their methodologies and tooling. While security researchers focus upon finding new and novel vulnerabilities, i.e., finding 0-days, security practitioners spend the majority of their time using known vulnerabilities and abusing security misconfigurations. These are often emulated through Capture-the-Flag (CTF) challenges which enable transfer learning. When attacking enterprise networks or performing privilege-escalation attacks, interviewees mentioned that they would never search for novel 0-day vulnerabilities due to their limited amount of time, but depend upon their knowledge of existing vulnerabilities as detailed by large online knowledge bases such as hacktricks [^6] https://book.hacktricks.xyz/.
@@
Linux Privilege-Escalation Vulnerabilities. Privilege-Escalation (short priv-esc) is the art of making a system perform operations that the current user should not be allowed to. We focus upon a subsection of priv-esc, namely local Linux low-privilege users trying to become the all powerful root system administrator indicated by an user id of 0. This is a common task that occurs after an initial system breach.
@@
Privilege-Escalation attacks are typically performed manually by searching for exploitable configurations or vulnerable tools. The initial act of system reconnaissance, often named enumeration, is often automated through usage of tools such as linpeas.sh [^7] https://github.com/peass-ng/PEASS-ng/tree/master/linPEAS. Exploitation itself is typically done manually through the, hopefully ethical, hacker.
@@
In many security areas, established standards and methodologies guide novice practitioners, e.g., in the web application area the non-profit organisation OWASP provides both the de-facto standard list of commonly used web vulnerabilities [^8] https://owasp.org/Top10/ as well as detailed testing guides [^9] https://owasp.org/www-project-web-security-testing-guide/stable/. In contrast, there is no such coverage for Linux Privilege Escalation Attacks. Partially fitting is the MITRE ATT&CK framework [^10] https://attack.mitre.org/resources/faq/#other-models-faq that “is a knowledge base of cyber adversary behavior and taxonomy for adversarial actions across their lifecycle”. Originally focusing upon Microsoft Windows Enterprise networks, subsequent iterations also include Linux attack vectors. MITRE ATT&CK does not offer a methodology, i.e., it does not describe attacks paths, but is an unordered taxonomy of potential attack vectors, thus does not provide high-level guidance to security practitioners, nor can be used as a high-level structure for benchmarks.
@@
Instead of established standards, aspiring penetration testers typically consume living online information sources. Ample unstructured information about Linux privilege escalation techniques can be found in public online wikis such as hacktricks [^11] https://book.hacktricks.xyz/ or GTFObins [^12] https://gtfobins.github.io/, a collection of privilege escalation techniques. In addition, Capture-the-Flag (CTF) style exercises allow penetration testers to hone their skills. Sites such as TryHackMe [^13] https://tryhackme.com/ or HackTheBox [^14] https://www.hackthebox.com/ allow online access to an ever-changing set of vulnerable virtual machines.
@@
Automated Linux Priv-Esc Tools. Kowira et al. (2024) give an overview of existing Linux enumeration scripts and state the lack of automated Linux privilege escalation. Penetration testers have to parse the various enumeration scripts’ outputs and match the provided information with potential exploitation attacks. In contrast, we investigate the usage of LLMs to autonomously enumerate and execute privilege-escalation attacks.
@@
Enumeration scripts such as linux-smart-enumeration [^15] Also often called lse.sh, https://github.com/diego-treitos/linux-smart-enumeration/tree/master, linPEAS [^16] https://github.com/carlospolop/PEASS-ng/tree/master/linPEAS or linenum.sh are rule-based. If paths are hard-coded, even simple obfuscation techniques, e.g., installing tools in different locations or running services on atypical ports, can avoid vulnerability detection. In addition, those tools lack situational awareness, i.e., they are not able to automatically integrate information within found documents, e.g., analyzing a stored email for saved passwords therein.

### v6

@@ Abstract
Penetration-testing is crucial for identifying and mitigating system vulnerabilities, with privilege-escalation being a critical subtask involving gaining elevated access to protected resources. The emergence of Large Language Models (LLMs) presents new avenues for automating these security practices by emulating human behavior. However, a comprehensive understanding of LLMs’ efficacy and limitations in performing autonomous Linux privilege-escalation attacks remains underexplored. To address this gap, we introduce hackingBuddyGPT, a fully automated LLM-driven prototype designed for evaluating autonomous Linux privilege-escalation. We curated a novel, publicly available Linux privilege-escalation benchmark comprising distinct, single-vulnerability virtual machines, enabling controlled and reproducible evaluation.
Our empirical analysis assesses the quantitative success rates and qualitative operational behaviors of various LLMs—GPT-3.5-Turbo, GPT-4-Turbo, and Llama3—against baselines of human professional penetration-testers and traditional automated tools. We investigate the impact of context management strategies, different context sizes, and various high-level guidance mechanisms on LLM performance.
Results show that GPT-4-Turbo demonstrates high efficacy, successfully exploiting 33–83% of vulnerabilities, a performance comparable to human penetration testers (75%). In contrast, local models like Llama3 exhibited limited success (0–33%), and GPT-3.5-Turbo achieved moderate rates (16–50%). High-level guidance significantly boosts LLM success rates, for instance when using GPT-4-Turbofrom 33% to 66% (without guidance) or from 66% to 83%, while state management through LLM-driven reflection doubled unaided GPT-4-Turbo success rates (from 33% to 66%).
Qualitative analysis reveals both LLMs’ strengths and weaknesses in generating valid commands and highlights challenges in common-sense reasoning, error handling, and multi-step exploitation, particularly with temporal dependencies. Cost analysis indicates that GPT-4-Turbo can achieve human-comparable performance at competitive costs per exploited vulnerability, especially with optimized context management. Our work provides a baseline for evaluating LLM capabilities in autonomous privilege escalation, guiding future research toward more effective and reliable LLM-guided penetration-testing.
@@ 1 Introduction
In the rapidly evolving field of cybersecurity, penetration-testing (“pen-testing” or “hacking”) plays a pivotal role in identifying and mitigating potential vulnerabilities. A crucial subtask of pen-testing is privilege-escalation, which involves exploiting a bug, design flaw, or configuration oversight in an operating system or software application to gain elevated access to resources that are normally protected from an application or user [^1] https://en.wikipedia.org/wiki/Privilege_escalation. The ability to escalate privileges provides a malicious actor with increased access, potentially leading to more significant breaches or system damage. Therefore, understanding and improving the performance of tools used for this task is highly relevant and impacts real-life security.
@@
In this paper, we focus on investigating the performance of Large Language Models (LLMs) in the context of penetration-testing, specifically for Linux privilege-escalation. LLMs have shown remarkable abilities to emulate human behavior that can be used to automate and improve various tasks in penetration-testing (Happe and Cito, 2023a; Deng et al., 2023). However, there is currently no understanding on how these models perform in common privilege-escalation scenarios. By understanding their performance, we can guide future research efforts towards higher effectiveness and reliability for LLM-guided penetration-testing, while ensuring the resulting tools are cost-effective and efficient for use by security practitioners operating under time constraints.
@@
To address this gap, we performed an empirical analysis of multiple LLMs using a newly created open-source Linux privilege-escalation benchmark, providing insight into LLMs’ strengths and weaknesses in the context of these attacks. We release a platform to evaluate and compare the performance of different LLMs in a controlled manner. By understanding the performance of these models in the critical task of privilege-escalation, we can guide future research efforts towards higher effectiveness and reliability for LLM-guided penetration-testing.
1.1 Motivation
In our previous work (Happe and Cito, 2023a), we employed a proof-of-concept autonomous hacking agent (wintermute) to attack a single vulnerable Linux virtual machine. Using GPT-3.5, we were able to experience successful privilege-escalation attacks occasionally. We will show in Section 2.3 that concurrent and subsequent research was able to confirm these offensive capabilities while additionally techniques such as Chain-of-Thought or Pentest-Task-Trees were able to improve results.
In this work, we want to investigate the latent knowledge and decision-making capabilities of off-the-shelf LLMs for Linux privilege-escalation attacks. This provides a baseline against which advanced techniques such as CoT can be compared against—if privilege-escalation attacks without these advanced techniques are already successful, their additionally needed resources can be saved.
1.2 Research Questions
We guide our work based on the following research questions:
  • RQ1: What is the efficacy of LLMs in performing autonomous Linux privilege-escalation attacks? This question includes multiple sub-questions:
    • How do the quantitative success rates and qualitative operational behaviors of autonomous LLM-based privilege-escalation agents compare against those of human penetration-testers and automated traditional privilege-escalation tools?
    • What are the primary challenging areas and qualitative limitations observed in LLM-generated commands?
  • RQ2: How do various context management strategies and context sizes impact the efficacy and efficiency of LLM-driven privilege-escalation agents?
  • RQ3: To what extent do different high-level guidance mechanisms influence the success rates of attack vectors by LLM-based privilege-escalation agents?
1.3 Contributions
@@
To answer our research questions (Section 1.2), we curated a Linux privilege-escalation benchmark, implemented an LLM-driven hacking prototype (hackingBuddyGPT), and identified properties of LLM-based penetration testing through empirical analysis. This approach results in the following contributions:
@@
  • a publicly available Linux privilege-escalation benchmark set that can be run on local premises due to the safety- and security-critical nature of this benchmark (Section 4).
@@
  • an fully-automated LLM-driven Linux privilege escalation-prototype (Section 3.3 HackingBuddyGPT: Autonomous Hacking Agent).
@@
We publicly release the source code of our prototype [^2] https://github.com/ipa-lab/hackingBuddyGPT, the created testbed [^3] https://github.com/ipa-lab/benchmark-privesc-linux, and captured trajectory data [^4] https://github.com/ipa-lab/hackingbuddy-results under an open-source license on github.
@@ 2 Background and Related Work
The background section focuses on the two distinct areas that this work integrates: LLMs and penetration-testing.
@@ 2.1 Large-Language Models
There is an ongoing discussion about the minimum viable model parameter size. On one hand, proponents claim that emergent features arise only with larger model sizes (Kosinski, 2023; Bubeck et al., 2023; Wei et al., 2022); on the other hand, proponents claim that smaller models can achieve domain-specific tasks with reduced costs for both training and execution (Bender et al., 2021).
Smaller models are feasible to run locally. This is important for agent-based scenarios (Andreas, 2022; Park et al., 2023) or if privacy reasons disallow the usage of cloud-based LLMs. In early 2024 the term Small Language Models was introduced to denote models with parameter sizes typically smaller than 8–12 billions, one example of such a model would be Llama3-8b.
An alternative to using small language models is quantizing models with larger parameter counts. In this approach, parts of the model weights are quantized from 32bit floating points into data types of lower precision, e.g., 4 bit integers. This reduces the model’s memory requirements, and thus makes local model usage computationally feasible. There is an ongoing discussion on the trade-off between using smaller models of full precision and using larger quantized models (Huang et al., 2024).
Training an LLM incurs high costs. Recently, alternative approaches have tried to achieve high performance while avoiding expensive training. In-Context Learning (ICL,  Dong et al. (2022); Dai et al. (2023)) includes background information within the prompt and thus exchanges knowledge inherently stored within the model with external knowledge. An alternative to ICL is Retrieval Augmented Generation (RAG, Lewis et al. (2020a)) in which the parametric memory of an LLM is extended by external non-parametric knowledge typically selected by a dedicated retrieval system. Recent research (Lee et al., 2024) shows that ICL rivals state-of-the-art RAG systems. Li et al. (2024) further indicate that ICL outperforms RAG in question-answering benchmarks while performing comparably for summarization tasks. These findings indicate that ICL can be used as a stand-in for RAG systems given that the used knowledge base fits into the LLM’s context size.
@@ 2.2 Penetration Testing
Penetration-Testing, short pen-testing, is described by Geer and Harthorne (2002) as “the art of finding an open door”. Its goal is to find a vulnerability within the subject-under-test to falsify the hypothesis that the subject is secure. The outcome of a penetration-test allows defenders to fortify their systems so that other, potentially malicious, attackers cannot abuse similar vulnerabilities (Bishop, 2007). Professionals performing those tests are typically called penetration-testers, pen-tester, or simply hackers. An additional differentiation is often performed upon the intend of the pen-tester: white-hats perform ethical research to improve the field of software security while black-hats are malicious and work for monetary or political gain.
@@
Shah and Mehtre (2015) further elaborate on the nature of penetration testing and differentiate between Vulnerability Assessments and Penetration Testing. The goal of the former is to identify as many possible vulnerabilities within the subject-under-test as possible, while the latter emulates an attacker that tries to actively exploit a found vulnerability. As penetration-testing can lead to system instabilities and data loss, automated tooling often focus upon Vulnerability Assessment, not exploitation (Shebli and Beheshti, 2018). Tooling such as nmap [^5] https://nmap.org/, OpenVAS [^6] https://www.openvas.org/, PortSwigger BURP [^7] https://portswigger.net/burp or ZAP [^8] https://www.zaproxy.org/ often utilize rule-based detection systems as well as databases of known vulnerable software versions and use aggressive techniques such as fuzzing only on explicit user interaction or as measures of last resort.
@@
Only little empirical research into how penetration-testers perform their work, and the potential problems therein, has been performed. Happe and Cito (2023b) performed an interview study with professional penetration-testers. One of their key findings is that security researchers and security practitioners (penetration-testers) differ in their methodologies and tooling. While security researchers focus upon finding new and novel vulnerabilities, i.e., finding 0-days, security practitioners spend the majority of their time using known vulnerabilities and abusing insecure configurations. These are often emulated through Capture-the-Flag (CTF) challenges indicating the possibility of transfer learning. When attacking enterprise networks, or performing privilege-escalation attacks, interviewees mentioned that they would never search for novel 0-day vulnerabilities due to their limited amount of time, further underscoring why efficiency and resource consumption (time, tokens, and monetary costs) are critical metrics in this domain. Instead they depend upon their knowledge of existing vulnerabilities as detailed by large online knowledge bases such as hacktricks (Polop, 2025).
2.2.1 Linux Privilege-Escalation Vulnerabilities
@@
Privilege-Escalation (short priv-esc) is the art of making a system perform operations that the current user should not be allowed to. We focus upon a subsection of priv-esc, namely local Linux low-privilege users trying to become the all powerful root system administrator indicated by an user id of 0. This is a common task that occurs after an initial system breach.
@@
Privilege-Escalation attacks are typically performed manually by searching for exploitable configurations or vulnerable tools. The initial act of system reconnaissance, often called enumeration, is automated through usage of enumeration tools such as linpeas.sh [^9] https://github.com/peass-ng/PEASS-ng/tree/master/linPEAS. These tools analyze the target system configuration and output a summary including potential avenues of attack. Exploitation itself is typically done manually through the, hopefully ethical, hacker.
@@
In many security areas, established standards and methodologies guide novice practitioners, e.g., in the web application area the non-profit organization OWASP provides both the de-facto standard list of commonly used web vulnerabilities (OWASP, 2021) as well as detailed testing guides (OWASP, 2013). In contrast, there is no such coverage for Linux Privilege Escalation Attacks. Partially fitting is the MITRE ATT&CK framework (Strom et al., 2018) that “is a knowledge base of cyber adversary behavior and taxonomy for adversarial actions across their lifecycle”. Originally focusing on Microsoft Windows Enterprise networks, subsequent iterations also include Linux attack vectors. MITRE ATT&CK does not offer a methodology, i.e., it does not describe attacks paths, but is an unordered taxonomy of potential attack vectors, thus does not provide high-level guidance to security practitioners, nor can it be used as a high-level structure for benchmarks.
@@
Instead of established standards, aspiring penetration testers typically consume living online information sources. Ample unstructured information about Linux privilege escalation techniques can be found in public online wikis such as hacktricks (Polop, 2025) or GTFObins (Pinna and Cardaci, 2025), a collection of privilege-escalation techniques. In addition, Capture-the-Flag (CTF) style exercises allow penetration testers to hone their skills. Sites such as TryHackMe [^10] https://tryhackme.com/ or HackTheBox [^11] https://www.hackthebox.com/ allow online access to an ever-changing set of vulnerable virtual machines.
2.2.2 Automated Linux Privilege-Escalation Tools
@@
Kowira et al. (2024) give an overview of existing Linux enumeration scripts and state the lack of automated Linux privilege-escalation. Penetration-testers have to parse the various enumeration scripts’ outputs and match the provided information with potential attacks. In contrast, we investigate the usage of LLMs to autonomously enumerate and execute privilege-escalation attacks.
@@
Enumeration tools such as linux-smart-enumeration [^12] Also often called lse.sh, https://github.com/diego-treitos/linux-smart-enumeration/tree/master, linPEAS [^13] https://github.com/carlospolop/PEASS-ng/tree/master/linPEAS or linenum.sh automate the often tedious tasks of gathering system information. They are rule-based: if paths are hard-coded, even simple obfuscation techniques, e.g., installing tools in different locations or running services on atypical ports, can avoid vulnerability detection. In addition, those tools lack situational awareness, i.e., they are not able to automatically integrate information within found documents, e.g., analyzing a stored email for saved passwords therein.
Automation in Linux privilege-escalation attack scenarios is typically focused on making system enumeration more efficient. Analysis of the gathered information as well as its exploitation is typically performed manually. This is a difference to the Windows-Ecosystem where attack tooling such as PowerUp.ps1 or SharpUp allows to both detect and automatically exploit insecure configurations.
The well known metasploit framework deprecated its automated exploitation module, autopwn_db in 2011 as “it did not fit in the scope of the framework, was unmaintained, and caused damage to systems when used in the default mode” (munky9001, 2011). Alternative solutions such as pwncat-cs [^14] https://github.com/calebstewart/pwncat or traitor [^15] https://github.com/liamg/traitor are infrequently updated.
2.2.3 Benchmarks and Testbeds